OpenStudy (anonymous):

How much money should be deposited today in an account that earns 9.5% compounded monthly so that it will accumulate to $10,000 in three years? (Do not round until the final answer. Then round to the nearest cent.)

OpenStudy (anonymous):

9.5% ??? where can i get in on that!!!

OpenStudy (blockcolder):

Let P be our mystery amount. Then P must satisfy the ff: \[10,000=P\left (1+\frac{0.095}{12}\right )^{3\cdot 12}\] Now, \[P=\frac{10,000}{\left (1+\frac{0.095}{12}\right )^{36}}\] Then plug this in a calculator. :D

OpenStudy (anonymous):

Thank you!